OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

y - k = m(x - h) where m is the slope of the line and (h, k) is a point on the line (any point works).

OpenStudy (anonymous):

Example: Write an equation of the line which passes through (3, 4) and has slope m = 5 . h = 3 and k = 4 . y - 4 = 5(x - 3) now can you plug in your info

OpenStudy (anonymous):

I got it, thanks.

OpenStudy (anonymous):

welcome